OpenStudy (anonymous):

please help (10x^2)/(6y^3) • (24y^2)/(35x^2)

OpenStudy (unklerhaukus):

\[\frac{10x^2}{6y^3} \cdot\frac{24y^2}{35x^2}\] \[=\frac{5\times2\times x^2}{6\times y^3} \cdot\frac{6\times 4\times y^2}{5\times 7\times x^2}\]

OpenStudy (unklerhaukus):

\[=\frac{\cancel5\times2\times \cancel {x^2}}{\cancel6\times y^{\cancel{3}}} \cdot\frac{\cancel6\times 4\times \cancel{y^2}}{\cancel5\times 7\times\cancel{ x^2}}\]\[=\frac{2\times4}{7\times y}\]

OpenStudy (anonymous):

thanks!! i see what i did wrong